sam is going to university this year. he wants to pay for his registration and buy himself a laptop so he can do his assignments and write his online exams. sam borrowed R15 000,00 from his grandfather at an interest rate of 9,5% per year, compounded monthly. Determine how much Sam will owe his grandfather after 5 years


1
Expert's answer
2021-06-01T10:35:56-0400

P= R15000

I = (9.5/12)%

= 19/24%

T =5*12

=60

A = p(1+r/100)t

=15000(1+(19/2400)60.

= 15000(1+0.0079)60

=15000(1.0079)60

=15000(1.605)


=R 24075.14